OpenStudy (anonymous):

Is this correct ? Express the radical in simplest form: sqrt 45x would be 9x sqrt 5

OpenStudy (eyust707):

\[\sqrt{45x}\]

OpenStudy (eyust707):

like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I think x should be in the sqrt(in the answer)

OpenStudy (anonymous):

\[3\sqrt{5x}\]

OpenStudy (eyust707):

\[\sqrt{45x} = \sqrt{9*5*x} = \sqrt{9}\sqrt{5x}\]

OpenStudy (eyust707):

yes!